The length of a rectangle is 8 more than 3 times the width, if the perimeter is 64, what are the measurements? The area?
Answer by stanbon(75887)

The length of a rectangle is 8 more than 3 times the width, if the perimeter is 64, what are the measurements? The area?
-----
L = 3W + 8
---
Perimeter = 2(length + width)
64 = 2(3W+8 + W)
32 = 4W+8
4W = 24
Width = 6
---
Then Length = 3*6 + 8 = 26
----
Area = L*W
Area = 6*26 = 156
=======================
Cheers,
Stan H.
==============